OpenStudy (anonymous):

In 1993 the radius of Hurricane Emily was about 350 km. The wind speed near the center ("eye") of the hurricane, whose radius was about 30 km, reached about 200 km/h. As air swirled in from the rim of the hurricane toward the eye, its angular momentum remained roughly constant. Estimate the wind speed at the rim of the hurricane?

OpenStudy (anonymous):

see the angular momentum is same for radius 350km, 30km angular momentum=m*v*r m=mass v=vel. r=rad. hence m*v*r=m*V*R m is cancelled 200km/h * 30km= V*350km V=17.14km/h